# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

### v1

@@ 3.3.1. Post-train Bayesian Strategy for Frequency-based Attack
(8) $\operatorname{logits}=g_{\theta^{\prime}}(f_{\theta}(\mathbf{x}))+f_{\theta_{c}}(\mathbf{x})$
@@
(9) $\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)}p\left(y\mid\mathbf{x},\Theta,\theta^{\prime}\right)\approx\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\mathbf{x},\Theta,\theta_{k}^{\prime}\right),\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)$
@@
(10) $\tilde{\mathbf{x}}^{i+1}=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}),\Theta,\theta^{\prime}_{k})\}$
3.3.2. Inference on Bayesian Appended Models
$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo(Springenberg et al., 2016) to sample appended model $\theta^{\prime}$ in each iteration:
$\theta^{\prime}_{t+1}=\theta^{\prime}_{t}-\sigma^{2}\mathbf{C}^{-1/2}_{\theta^{\prime}_{t}}\mathbf{h}_{\theta^{\prime}_{t}}+\mathbf{N}(0,2F\sigma^{3}\mathbf{C}^{-1}_{\theta^{\prime}_{t}}-\sigma^{4}\mathbf{I})$
(11) $\mathbf{C}_{\theta^{\prime}_{t}}\leftarrow(1-\tau^{-1})\mathbf{C}_{\theta^{\prime}_{t}}+\tau^{-1}\mathbf{h}_{\theta^{\prime}_{t}}^{2}$
where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically (Springenberg et al., 2016).
3.4. Hybrid Adversarial Attack
Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain (Wang et al., 2020; Jeong et al., 2022). We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains. Specifically, we define the hybrid attack as:
(12) $\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}(g^{i}_{k}+d^{i}_{k})\}$
(13) $g^{i}_{k}$ $=\frac{1}{N}\sum_{n=1}^{N}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}_{n-1}),\Theta,\theta^{\prime}_{k}),\tilde{\mathbf{x}}^{i}_{0}=\tilde{\mathbf{x}}^{i}$
(14) $d^{i}_{k}$ $=\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta,\theta^{\prime}_{k})$
@@ 3.4. Hybrid Adversarial Attack
Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$;$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models; ;
@@
                   Update $\theta^{\prime}_{k}$ with $\mathbf{h}_{\theta^{\prime}_{k}}$ via Sec. 3.3.2;
@@
             Averaging frequency gradient $g_{k}$ using Eq. 13 ;
             Calculating spatial gradient $d_{k}$ using Eq. 14 ;
@@
4. Experimets
4.1. Experimental Settings
@@
Datasets: We choose three generated image datasets created by a wide range of generative models, across AIGI detection and Deepfake detection. Synthetic LSUN is a commonly used dataset proposed by CNNSpot (Wang et al., 2020), containing 720k real images from LSUN and 360k fake images generated by ProGAN (Karras et al., 2018). GenImage (Zhu et al., 2024) is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models. Following the protocol in (Zhu et al., 2024), we employ a subset of GenImage collecting 162k real images from Imagenet (Deng et al., 2009) and 162k Stable Diffusion(SD) V1.4 (Rombach et al., 2022) generated images for training. The images generated by the other generators are used for testing in Sec. 4.3.2. To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by (Shamshad et al., 2023). The Deepfake dataset consists of 50k real face images from FFHQ (Karras et al., 2019) and 50k generated face images generated by StyleGAN2 (Karras et al., 2019). After training, we collect only the correctly classified testing samples for attack in evaluation.
@@
Evaluated Models: We extensively evaluate the transferability of adversarial examples by eight different models across heterogeneous architectures under the spatial and frequency domain. We choose spatial-based classification models ResNet-50 (He et al., 2016) and MobileNet (Howard et al., 2017) as the source models. For target models, apart from the above models, we also use spatial-based models including EfficientNet (Tan and Le, 2019), Vision Transformer(Vit) (Dosovitskiy et al., 2020) and Swin-Transformer(Swin-Vit) (Liu et al., 2021), and frequency-based detectors DCTA (Frank et al., 2020) and Spec (Zhang et al., 2019). We report the accuracy of these detectors in the supplement. Besides, we also evaluate the robustness of state-of-the-art AIGI detection methods, including GramNet (Liu et al., 2020), LGrad (Tan et al., 2023), LNP (Liu et al., 2022) and UnivFD (Ojha et al., 2023).
@@
Compared Methods: We adopt gradient-based methods, I-FGSM (Kurakin et al., 2018), PGD (Madry et al., 2017), MI (Dong et al., 2018) and Frequency-based methods S2(Long et al., 2022). We also consider ensemble-based methods Ensemble (Dong et al., 2018) and SVRE (Xiong et al., 2022). For a fair comparison, we ran 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all attacks.
@@
Implementation Details: We choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entroy loss with an initial learning rate of 0.0001. We follow the same data augmentation strategy in CNNSpot (Wang et al., 2020). Before cropping, images are blurred with $\sigma$ Uniform[0, 3] with 10% probability, and JPEG-ed with 10% probability. We transformed images to 224 pixels and apply ImageNet normalization. For post-train Bayesian optimization, we follow the default setting in (Wang et al., 2023b). Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate. Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$. For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following (Long et al., 2022). All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
@@
Table 1. The attack success rate(%) on CNN-based, Vit-based and Frequency-based models on the Synthetic LSUN, GenImage subset and Synthetic FFHQ datasets. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
@@
Surrogate Model Attack Methods ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Average
@@
ResNet PGD 78.3 77.6 49.4 73 25.6 47.6 41.0 70.2 54.9
S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48 51.5 94.8 69.2
@@
IFGSM 67.5 67.5 55.3 67.5 28.9 32.5 16.2 17.5 40.8
MIFGSM 67.5 67.5 54.9 67.6 28.8 32.6 20.7 17.9 41.4
ResNet PGD 100 100 88.6 99.9 49.9 50 40.7 50.0 68.4
S2I 100 100 77.7 90.1 46.7 52.9 37.6 55.8 65.8
FPBA(Ours) 100 100 86.8 100 49.4 49.6 43.1 50.9 68.5
IFGSM 10.7 12.7 23.8 60.4 24.5 28.4 7.3 10.2 16.8
MIFGSM 10.3 12.6 23.2 60.4 25.4 28.7 9.1 10.2 17.1
MobileNet PGD 49.8 50.2 57.5 100 49.7 49.8 38.8 49.8 49.4
S2I 82.8 86.7 78.5 100 51.4 51.8 28.9 49.7 61.4
FFHQ (Style GAN) FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
4.2. The Evaluation on Normally Trained Models
We show the attack performance against all models on all datasets in Tab. 1. Under the white-box setting, our proposed method FPBA achieves the highest attack success rate in all cases, and outperforms other competitive methods. Specifically, FPBA gets an average white-box success rate across different datasets and models as high as 99.8%, while S2I, PGD, IFGSM and MIFGSM only have 95.6%, 95.4%, 65.4% and 65.5%. Under the black-box setting, FPBA still achieves the highest average transfer success rate of 55.3%, surpassing MIFGSM, PGD and S2I by a margin of 27.7%, 7.6% and 8.4% respectively.
Next, We find that the attack results for FPBA are usually not the best on frequency-based models. To further investigate the reason, we plot the spectrum saliency map of frequency-based detectors using Eq. 4 in Fig. 5. Unlike spatial-based detectors, which rely on numerous frequency components to make decisions (see Fig. 3), the frequency components focused on in frequency-based models are very sparse. Therefore, the gradient information available from the spatial domain is much richer compared to the frequency domain. Nevertheless, our method still has the best average success rate over all models. In addition, we combine spatial-based gradient and frequency-based to further improve the attack performance on frequency-based models. FPBA gets an average transfer success rate of 43.1% against frequency-based detectors, which is only slightly lower 1.5% than PGD, and much higher 3.3%, 15.8% and 17.6% than SSA, MIFGSM and IFGSM respectively.
Finally, FPBA has nearly top-2 results in all 48 cases (datasets vs. victim models vs. target models vs. attack setting), which shows consistent attack performance across different attack settings (white-box vs. black-box), tasks (AIGI detection vs. Deepfake detection) and models (CNN-based vs. Vit-based vs. Frequency-based detectors). We leave the detailed robustness analysis across models to Sec. 4.4.1.
Figure 5. Visualization of the sensitive frequency components of real and fake images (average 1000 images on LSUN/ProGan datasets) for frequency-based models. The frequency components of frequency-based models are highly sparse in comparison with spatial-based models.
Table 2. The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets. 1 means setting (1), i.e. taking an ensemble of ResNet-50, MobileNet, and EfficientNet; 2 means setting (2), i.e. taking an ensemble of ResNet, MobileNet and ViT; 3 means setting (3), i.e. taking an ensemble of ResNet, DCTA and ViT. 4 means only using ResNet as surrogate model.
Attack ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave
@@ 4.2. The Evaluation on Normally Trained Models
4.2.1. Comparsion with Ensemble-based Attacks
Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods, including ENSEMBLE  (Dong et al., 2018) and SVRE (Xiong et al., 2022). Since there are three types of models for AIGI detection (CNN-based vs. Vit-based vs. Frequency-based detectors), we hence conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (ResNet-50, MobileNet, EfficientNet); (2) taking CNNs and Vits as ensemble surrogates (ResNet-50, MobileNet, Vit); (3) taking CNNs, Vits and frequency-based detectors as ensemble surrogates (ResNet-50, DCTA, Vit). Although we can also use more than one architecture for our method, we only use ResNet-50 as the surrogate architecture to verify the universal transferability across heterogeneous models. We report the results in Tab. 2. First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across Vit-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors. We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate. Second, FPBA solely on ResNet still achieves competitive results in comparison with ensemble methods. Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases. For instance, the adversarial example only generated by ResNet-50 with FPBA get success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of ResNet-50, EfficientNet and MobileNet. This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
Table 3. Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot (Wang et al., 2020).
Attack GramNet LGrad LNP UnivFD Ave
IFGSM 39.4 7.5 37.0 14.6 24.6
MIFGSM 9.65 6.6 38.5 24.2 19.7
PGD 67.7 50.2 47.4 11.5 44.2
S2I 49.1 49.1 28.3 8.3 33.7
FPBA 87.8 50.2 39.5 16.6 48.5
4.3. The Evaluation on Diverse AIGI Detections
4.3.1. Attack SOTA AIGI Detectors
In this section, we examine the robustness of state-of-the-art (SOTA) AIGI detectors against adversarial attacks and present the results in Tab. 3. Even though these detectors can achieve SOTA detection performance on benign data, they still suffer from a serious adversarial threat and demonstrate our method is the strongest threat to these SOTA detectors. Unlike common detectors trained on the fake images for the task of binary classification, UnivFD (Ojha et al., 2023) extract fixed features via CLIP (Radford et al., 2021) and use nearest neighbor or linear probing for classification. The obvious diversity of the surrogate model(CNNSpot) and victim model(UnivFD) leads to a limited adversarial transferability on each other. Nevertheless, our method still achieves the best results on UnivFD compared to other attacks.
Table 4. Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
@@ 4.3.1. Attack SOTA AIGI Detectors
Table 5. The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-Vit.

### v3

@@ III-C1 Post-train Bayesian Strategy
$\operatorname{logits}=g_{\theta^{\prime}}(f_{\theta}(\mathbf{x}))+f_{\theta_{c}}(\mathbf{x}).$ (7)
@@
$\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)}p\left(y\mid\mathbf{x},\Theta,\theta^{\prime}\right)\approx\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\mathbf{x},\Theta,\theta_{k}^{\prime}\right),$
$\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta),$ (8)
@@
$\tilde{\mathbf{x}}^{i+1}=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}),\Theta,\theta^{\prime}_{k})\}.$ (9)
III-C2 Inference on Bayesian Appended Models
$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo[42] to sample appended model $\theta^{\prime}$ in each iteration:
$\theta^{\prime}_{t+1}=\theta^{\prime}_{t}-\sigma^{2}\mathbf{C}^{-1/2}_{\theta^{\prime}_{t}}\mathbf{h}_{\theta^{\prime}_{t}}+\mathbf{N}(0,2F\sigma^{3}\mathbf{C}^{-1}_{\theta^{\prime}_{t}}-\sigma^{4}\mathbf{I}),$
$\mathbf{C}_{\theta^{\prime}_{t}}\leftarrow(1-\tau^{-1})\mathbf{C}_{\theta^{\prime}_{t}}+\tau^{-1}\mathbf{h}_{\theta^{\prime}_{t}}^{2},$ (10)
where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically [42].
III-D Hybrid Adversarial Attack
Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain [3, 21]. We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains. Specifically, we define the hybrid attack as:
$\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}(g^{i}_{k}+d^{i}_{k})\},$ (11)
$g^{i}_{k}$ $=\frac{1}{N}\sum_{n=1}^{N}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}_{n-1}),\Theta,\theta^{\prime}_{k}),\tilde{\mathbf{x}}^{i}_{0}=\tilde{\mathbf{x}}^{i},$ (12)
$d^{i}_{k}$ $=\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta,\theta^{\prime}_{k}),$ (13)
@@ III-D Hybrid Adversarial Attack
Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$;$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models;
@@
                   Update $\theta^{\prime}_{k}$ with $\mathbf{h}_{\theta^{\prime}_{k}}$ via Section III-C2;
@@
             Average frequency gradient $g_{k}$ using Eq. 12 ;
             Calculate spatial gradient $d_{k}$ using Eq. 13 ;
@@
IV Experimets
IV-A Experimental Settings
@@
Datasets: We choose three generated image datasets created by a wide range of generative models, across AIGI detection and Deepfake detection. Synthetic LSUN is a commonly used dataset proposed by CNNSpot [3], containing 360k real images from LSUN and 360k fake images generated by ProGAN [43]. GenImage [12] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models. Following the protocol in [12], we employ a subset of GenImage collecting 162k real images from Imagenet [44] and 162k Stable Diffusion(SD) V1.4 [45] generated images for training. The images generated by the other generators are used for testing in Section IV-C2. To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [46]. The Deepfake dataset consists of 50k real face images from FFHQ [47] and 50k generated face images generated by StyleGAN2 [47]. After training, we collect only the correctly classified testing samples for attack in evaluation.
@@
Evaluated Models: We extensively evaluate the transferability of adversarial examples by 13 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods. For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [3], MobileNet [48], EfficientNet [49] and DenseNet [50], and ViT-based detectors Vision Transformer(ViT) [51] and Swin-Transformer(Swin-ViT) [52]. For evaluating on various detection methods, we use frequency-based detectors DCTA [11] and Spec [18], gradient-based detectors LGrad [23], CLIP-based detectors UnivFD [24], and diffusion-based detectors DNF [28]. In addition, LNP [22] extracts the noise pattern of images and GramNet [53] learns the global texture representation. We also consider them as victim models.
@@
Compared Methods: We adopt gradient-based methods, I-FGSM [33], PGD [38], MI [54] and frequency-based methods S2[39], SSAH [55] and ensemble-based methods ENS [54] and SVRE [56]. Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks Fakepolisher [36] and TraceEvader [37] as baselines. Fakepolisher and TraceEvader remove the Deepfake traces and we follow their default settings. For other iterative attacks, we run 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all these attacks.
@@
Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [3]. Specificaly, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entroy loss with an initial learning rate of 0.0001. For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [3] to improve the models’ generalization and robustness. Before cropping, images are blurred with $\sigma\sim\text{Uniform}[0,3]$ with 10% probability, and JPEG-ed with 10% probability. We transformed images to 224 pixels on LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [3]. On the FFHQ (StyleGAN2) dataset we resized images to 224 pixels to ensure the integrity of the real/fake faces. Subsequently, we apply ImageNet normalization across three datasets. For other AIGI detectors, we use the pre-trained model from their official code. For post-train Bayesian optimization, we follow the default setting in [57]. Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate. Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$. For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [39]. All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
@@
TABLE I: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic LSUN and GenImage subset. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model. We mark the white-box attack results in gray, and black-box attack results are not marked with colors.
@@
Surrogate Model Attack Methods CNNSpot DenseNet EfficientNet MobileNet Spec DCTA ViT Swin Average
@@
PGD 78.3 77.6 49.4 73.0 25.6 47.6 41.0 70.2 54.9
CNNSpot S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
SSAH 97.8 2.5 1.3 3.0 0.8 1.0 1.0 0.7 1.4
FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
@@
IV-B Attack on Spatial-based and Frequency-based Detectors
We report the attack performance against spatial-based and frequency-based detectors in Table I. Under the white-box setting, our proposed method FPBA achieves the highest attack success rate in all cases, and outperforms other competitive methods. Specifically, FPBA gets an average white-box success rate across different datasets and models as high as 99.6%, while S2I, PGD, IFGSM and MIFGSM only have 93.1%, 93.1%, 66.2% and 66.2%. Although SSAH has a relatively high white-box success rate (97.7%), it fails to transfer the adversarial examples to black-box models. Under the black-box setting, FPBA still achieves the highest average transfer success rate of 50.5%, surpassing the IFGSM, MIFGSM, PGD, S2I and SSAH by a big margin of 25.7%, 23.3%, 8.4%, 11.9% and 49.1%. Note that our proposed FPBA significantly outperforms the SOTA frequency-based attack S2I and SSAH, demonstrating that FPBA is a stronger frequency-based attack against AIGI detectors.
TABLE II: The attack success rate(%) compared with attacks agaisnt Deepfake detectors(only fake ASR).DenNet, EffNet, MobNet refers to DenseNet, EfficientNet, MobileNet respectively.
Datasets Attack CNNSpot DenNet EffNet MobNet ViT Swin Ave.
FakePolisher 94.5 98.3 71.6 88.6 15.1 99.5 77.9
LSUN TraceEvader 0.6 1.3 4.5 4.8 8.10 11.4 1.9
(ProGAN) FPBA 100.0 100.0 98.2 99.5 38.6 99.9 89.4
FakePolisher 93.5 93.2 12.2 95.5 33.5 97.6 70.9
GenImage TraceEvader 46.1 34.3 9.4 37.0 15.2 44.3 31.1
(SD) FPBA 100.0 99.7 97.7 98.9 42.1 97.9 89.4
Figure 5: Visualization of the sensitive frequency components of real and fake images (average 1000 images on LSUN/ProGAN datasets) for frequency-based models. The frequency components of frequency-based models are highly sparse in comparison with spatial-based models. The color value represents the absolute gradient value of the model loss function after max-min normalization.
IV-B1 Comparison with Attacks against Deepfake Detectors
TABLE III: The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets. 1 means setting (1), i.e. taking an ensemble of CNNSpot, MobileNet, and EfficientNet; 2 means setting (2), i.e. taking an ensemble of CNNSpot, MobileNet and ViT; 3 means setting (3), i.e. taking an ensemble of CNNSpot, DCTA and ViT. 4 means only using CNNSpot as surrogate model.
Attack CNNSpot DenNet EffNet MobNet Spec DCTA ViT Swin Ave
@@ IV-B1 Comparison with Attacks against Deepfake Detectors
TABLE IV: Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [3].
Attack GramNet LGrad LNP UnivFD DNF Ave
MIFGSM 9.7 6.6 38.5 24.2 46.7 25.1
PGD 67.7 50.2 47.4 11.5 48.3 45.0
S2I 49.1 49.1 28.3 8.3 43.0 35.6
FPBA 87.8 50.2 39.5 16.6 49.8 48.9
TABLE V: Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
@@
Attacks against Deepfake detectors are similar to our attacks. However, Deepfake mainly manipulating or synthesizing faces, often crafted by GANs. In contrast, AIGI detectors identify any type of visual content generated by diverse techniques, such as GANs, diffusion models, and autoregressive models. The content includes humans, non-human entities, abstract art, landscapes, and more. These differences make it challenging to transfer attacks from DeepFake detectors to AIGI detectors. For example, techniques like searching for adversarial points on the face manifold [35] or removing facial forgery traces [36, 37] may not be applicable for AIGI detectors. To demonstrate this, we compare with FakePolisher [36] and TraceEvader [37], which are both the SOTA attacks for Deepfake detection. Because FakePolisher and TraceEvader are designed for removing forgery traces, we follow their default setting to only report the attack success rate on fake images. As shown in Table II, FPBA significantly outperforms FakePolisher and TraceEvader on both AIGI datasets. Although FakePolisher performs better than TraceEvader, but the former largely reduces the image quality, which is very visible and raises suspicion (Fig. 7). We suspect this is because FakePolisher has only learned the face manifold, limiting its ability to reconstruct fine-grained details in AIGI images.
IV-B2 Comparison with Ensemble-based Attacks
Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [54] and SVRE [56], i.e. utilizing an ensemble of surrogate models to generate adversarial examples. To examine the impact of different surrogate combinations(CNN-based vs. Vit-based vs. Frequency-based detectors), we conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (CNNSpot, MobileNet, EfficientNet); (2) taking CNNs and ViTs as ensemble surrogates (CNNSpot, MobileNet, ViT); (3) taking CNNs, ViTs and frequency-based detectors as ensemble surrogates (CNNSpot, DCTA, ViT). Although we can also use more than one architecture for our method, we only use CNNSpot as the surrogate architecture to verify the universal transferability across heterogeneous models. We report the results in Table III. First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across ViT-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors. We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate. Second, FPBA solely on CNNSpot still achieves competitive results in comparison with ensemble methods. Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases. For instance, the adversarial example only generated by CNNSpot with FPBA get success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of CNNSpot, EfficientNet and MobileNet. This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
IV-B3 Evaluation on Frequency-based Detectors
Although FPBA has the best average success rate, we find that the attack results are usually not the best on frequency-based models. To further investigate the reason, we plot the spectrum saliency map of frequency-based detectors using Eq. 3 in Fig. 5. Unlike spatial-based detectors, which rely on numerous frequency components to make decisions (see Fig. 3), the frequency components focused on in frequency-based models are very sparse. Therefore, the gradient information available from the spatial domain is much richer compared to the frequency domain. Nevertheless, FPBA still gets an top-2 average transfer success rate of 43.1% against frequency-based detectors, which is only slightly lower 1.5% than PGD, and much higher 3.3%, 15.8% and 17.6% than S2I, MIFGSM and IFGSM respectively.
TABLE VI: The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-ViT.